SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Jan.3, 2006)

Attendees: Al Eisner, Fang Fang, Selina Li, Dong Su.

Operations

  • Overall, Trigger has been running smoothly for the last two weeks over the winter break.
  • The old TSF-Y4 had a small problem in run 60597-60600 during the owl shift of Dec.26, 2005. No impact on the operations or data quality. See HN posted.
  • The TSF-Y crate had to be powered back on due to a power glitch on Dec.27, 2005 owl shift. Some other sub-detectors had similar problem too.
  • There will be ROD tomorrow (Wednesday) and will last about 1-1.5 shifts. This is an opportunity to do some tests at IR2.

DCZ

  • We will test the TSF-X version 9 of the firmware at IR2 tomorrow morning.

Online

  • Fang will test the DCZ online lib in which the test stand dependency was removed at IR2 tomorrow morning. Last time we saw large deadtime due to DCT/DCZ when it was tested on Dec.21, 2005.

Data Quality and Monitoring

  • The monitoring of loss of trickle injection was briefly discussed. See the relevant discussions in the detector operaton HN posted. Al mentioned that Rainer might be able to help or provide some insights.

Run 5 "startup" tasks


This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.